OpenStudy (anonymous):

Is it possible that the car could have accelerated to 55mph within 268 meters if the car can only accelerate from 0 to 60 mph in 15 seconds?

OpenStudy (anonymous):

find the time by using t= x/v

OpenStudy (anonymous):

u have to convert the units since u have miles per hour and seconds

OpenStudy (amistre64):

vf^2 = vi^2 + 2ax sounds familiar to me 55^2 = 2a (268 meters converted to miles) unless we are doing meters per hour?

OpenStudy (phi):

can you change 60 miles per hour to meters per second ? 1 mile = 1609.3 meters